OpenStudy (anonymous):

help plz What is the volume of the pyramid? A. 120 ft3 B. 200 ft3 C. 300 ft3 D. 600 ft3

OpenStudy (anonymous):

Guessing A.

OpenStudy (anonymous):

here is the question

OpenStudy (anonymous):

anyone? no? ok .-. i knew it

Vocaloid (vocaloid):

volume = (1/3)*base*(height) the base is a square with sides 5 and 5, so what is the area of the square?

OpenStudy (anonymous):

25

Vocaloid (vocaloid):

good, now multiply that by the height

OpenStudy (anonymous):

20*

Vocaloid (vocaloid):

no, it's 25 not 20

OpenStudy (anonymous):

4 * 5 is 20

OpenStudy (anonymous):

oh ok xD

Vocaloid (vocaloid):

area of a square = side*side = 5*5

OpenStudy (anonymous):

ok

Vocaloid (vocaloid):

now multiply that by the height

OpenStudy (anonymous):

25 x 25?

Vocaloid (vocaloid):

the height is 24 so 25*24 = ?

OpenStudy (anonymous):

25 x 24

Vocaloid (vocaloid):

right, and what does that equal?

OpenStudy (anonymous):

600

OpenStudy (anonymous):

D

Vocaloid (vocaloid):

nope we're not done yet

Vocaloid (vocaloid):

now divide that by 3 and you'll have your answer..

OpenStudy (anonymous):

200

Vocaloid (vocaloid):

good

OpenStudy (anonymous):

yay! thank you!